# Operations & Supply Chain Management: Principles and Practice

### v1

@@ Conclusion
Driven by technological progress, like in the fields of sensors, real-time control, autonomous driving, and robotisation, AMRs have become more and more successful in recent years (Fragapane et al., 2021). Other than their famous predecessors, Automated Guided Vehicles (AGVs), AMRs are not bound to predefined paths; computer vision allows flexible adaptation to a changing environment. In picker-to-parts systems, AMRs that also pick, such as the Toru robot of Magazino, can substitute human pickers, at least for a restricted product assortment (e.g., shoeboxes Boysen and De Koster, 2025). For non-standard products, transport AMRs accompany human pickers through the warehouse and relieve them of the burden of transporting the picked products back to the depot (Azadeh et al., 2023; Löffler et al., 2022). The most famous AMRs of parts-to-picker systems are (probably) shelf-lifting (KIVA) robots. They drive under a rack (also denoted as inventory pod), lift it, and deliver the pod to a stationary picker (Weidinger et al., 2018). Other AMR representatives are storage grid-supported (Autostore) robots (Azadeh et al., 2019), rack-climbing (Exotec) robots (Chen et al., 2022), and multi-tote (Quicktron, Haipick) AMRs (Qin et al., 2024) delivering bins autonomously from storage locations. Finally, there are also AMRs for order consolidation (Zou et al., 2021). Equipped with a tiltable tray or small conveyor tablet, they can distribute goods between sorting destinations.
@@
This section explores the concepts, mathematical modelling, and recent advancements in solving the BAP. For a more comprehensive description of the topic, see the latest review articles, such as Demir et al. (2014) and Weiss et al. (2019).
@@
The BAP has been formulated in three different ways according to the objective function and the constraints. The machines are given and the buffer sizes $N_{i}$ are found.
Problem Formulation. The goal of Gershwin and Schor (2000) was to minimise the total buffer size subject to a production rate constraint:
@@
where $\bar{n}_{i}$ is the average inventory level of the $i$th buffer, and $\mathbb{E}[LT(\mathbf{N})]$ and $\mathbb{E}[WT(\mathbf{N})]$ represent the average lead time and average waiting time of customer orders. Yuzukirmizi and Smith (2008) include the lead time in their objective function. Shi and Gershwin (2009) solve the profit maximisation problem subject to the nonlinear production rate constraint.
@@
For long lines, the generalised expansion method (Kerbachea and Smith, 1987) is proposed for systems with reliable machines, while decomposition and aggregation are prevailing methods to deal with unreliable machines. The decomposition method (Gershwin, 1987; Dallery et al., 1988), and its extensions (e.g., Tolio and Matta, 1998; Levantesi et al., 2003) enabled the evaluation of different models of long lines with unreliable machines. It approximates a $K$-machine $K-1$-buffer system with $K-1$ two-machine one-buffer systems. Decomposition aims to find parameters for these two-machine systems so that, for each buffer, the material flow behaviour in the two-machine line closely matches that in the original line. The goal is for an observer of the flow of parts into and out of a buffer in a real production line can be convinced that he is observing the flow of parts in a fictitious two-machine line. The aggregation method (Lim et al., 1990) combines all the two-machine one-buffer subsystems into a single aggregated machine. It consists of a forward aggregation which merges the first subline from the start towards the end and a backward aggregation which merges the last subline into a single machine from the end towards the beginning. This is repeated until the throughput converges. Simulation is another tool for the performance evaluation of large-scale production systems. It is flexible and can be accurate but it is time-consuming (Helber et al., 2011).
Optimisation techniques. Gradient methods use approximate techniques such as decomposition to calculate the gradient via forward difference. A line search is then conducted in that direction until a maximum is encountered and this becomes the next guess (Shi and Gershwin, 2009). Although buffer sizes $N_{i}$s are integers, the analytical form of the two-machine line evaluation treats each $N_{i}$ as a continuous variable, and therefore a gradient can be calculated. After finding the optimal continuous solution, buffer sizes can be rounded to the nearest integers.
Meta-heuristic methods treat buffer sizes as integers. Examples include genetic algorithms (Dolgui et al., 2002), tabu search (Demir et al., 2011), simulated annealing (Spinellis et al., 2000), and degraded ceiling (Nahas et al., 2006). Hybrid algorithms that combine multiple meta-heuristics include Shi and Men (2003) and Kose and Kilincci (2015). Mathematical programming techniques include the dynamic programming which breaks the problem into smaller subproblems (Diamantidis and Papadopoulos, 2004). Colledani and Tolio (2005) linearise decomposition equations, converting the problem into a mixed-integer linear programming problem.
@@
In modern manufacturing systems, the decreasing product lifetimes necessitate frequent factory reconfigurations or replacements, limiting the time available to learn and improve these systems. As a result, intuition derived from research and from careful, systematic observation of real factories becomes increasingly vital (Gershwin, 2017). Academic research provides theoretical insights and models and offers valuable tools and methodologies that can be directly applied to improve industrial processes. Works such as Burman et al. (1998), Liberopoulos and Tsarouhas (2002), Patchong et al. (2003), Colledani et al. (2010), Matta et al. (2012), and Li (2013) demonstrate how theoretical advancements can be translated into effective solutions for real-world challenges.
@@
Traditionally, production control focuses on internal information flows and can generally be categorised as push (e.g., make-to-stock, driven by demand forecast) or pull (e.g., make-to-order, triggered by customer demand; Lödding, 2011; Thürer et al., 2024). With push control, the flow of materials is pushed through successive processes from suppliers to inventory and then queues, workstations, buffers, etc. until the product is complete and ready to be delivered to the customer. In pull systems, production is triggered by an order (customer demand) which pulls the materials through the processes with the aim to minimise inventory and overproduction. In modern production control systems, hybrid systems are often used to combine push and pull logic. Common hybrid approaches include drum-buffer-rope and constant work in progress (CONWIP; Spearman et al., 1990). Alternative Material Flow Control (MFC) methods have been proposed, such as order generation, order release and production authorisation (Thürer et al., 2024). The dominant trend remains with pull approaches as Lean, Kanban and Just-In-Time (JIT) systems have become established best practices for cost and waste reduction. This is further reinforced with the increased integration of external information flows with the rise of customisation, penalisation and engineer-to-order models.
@@
With Industry 4.0 developments and implementations, increased digitisation and Industrial Internet of Things (IIoT) are becoming more prevalent (Kolberg and Zühlke, 2015). However, the amount and richness of data can be overwhelming, especially for SMEs (Moeuf et al., 2018). Machine learning and predictive methods used in simulation and optimisation models are promising ways to use the data for decision making and to increase production efficiency, flexibility and resilience (Martins et al., 2020). Digital twins for real-time monitoring and closed-loop feedback systems can address the challenges of adapting to rapidly changing manufacturing environments. The ISO 23247 standardisation effort on ‘Automation systems and integration – Digital twin framework for manufacturing’ are beneficial for structuring production planning and control systems (International Organization for Standardization, 2021; Shao et al., 2023), including, for example, Part 1: Overview and general principles, Part 2: Reference architecture, Part 3: Digital representation of manufacturing elements, and Part 4: Information exchange. Additional validation and more detailed specification work including validation with demonstrators are ongoing as of now. Blockchain-like technologies are also on the rise to facilitate multi-stakeholder interoperability across value chains to enable traceability and transparency in production processes and along these value chains.
@@
  • Single-model lines: A single homogeneous product is manufactured on the line (Otto et al., 2013). During each cycle time, every workstation performs the same set of tasks. Consequently, the entire set of tasks must be divided among the workstations, with each workstation handling a specific subset.
@@
Scheduling is the allocation of resources to tasks over time. The result is a schedule, which gives information on the start time and the completion time of each task as well as the times when resources are processing (which) tasks throughout the planning horizon. Synonyms for tasks in the literature are jobs or activities. Scheduling is one of the major and classical planning problems in operations management, see, e.g., §2.4, §2.11, §2.16, §2.18. Recently, scheduling is also addressed in supply chain management, see Amaro and Barbosa-Póvoa (2008), Chen and Hall (2022), Kreipl and Pinedo (2004), and Sawik (2011). In most cases, scheduling is a short- to mid-term planning problem, where the planning horizon stretches from a few hours to several months. Scheduling is done in manufacturing and services, see, e.g., Pinedo (2009). In manufacturing, tasks relate to physical products, which have to be processed in production facilities. In services, the main fields for scheduling problems are in transportation (e.g., trucks, which have to (un)loaded at terminal gates, see Boysen and Fliedner, 2010, ships which have to be (un)loaded at berths, see Bierwirth and Meisel, 2010, and Bierwirth and Meisel, 2015, containers, which have to be moved by cranes, see Boysen et al., 2017, aircrafts, which have to use runways for starting and landing, see Ikli et al., 2021), healthcare (e.g, patients, which have to be treated by doctors, see Hall, 2012, and Youn et al., 2022), sports (see Kendall et al., 2010; Ribeiro, 2013)), and workforce scheduling (see Van den Bergh et al., 2013). See also Pinedo (2009) and Pinedo et al. (2015) for a survey of scheduling in services.
Tasks
While being processed a task occupies resources. In the most simple case the resource is occupied completely and cannot handle any other tasks in parallel. However, tasks also might seize a fraction of a resource’s capacity and, if it is sufficiently small, the resource might handle other tasks in parallel. A task’s processing time specifies how long it occupies the resource. In most cases, the processing time is a given parameter. In more involved settings the decision maker might have the opportunity to decide on the processing time (Weglarz et al., 2011; Shabtay and Steiner, 2007; Gawiejnowicz, 2020). Once processing of a task has started it may or may not be preempted. However, processing a task is only allowed within a time interval between its release date and its deadline. In (quite common) special cases, release dates and deadlines may not impose any relevant restrictions and we can think of them as non-existent.
Precedence constraints
The times when tasks are processed are interdependent through the joint need for scarce resources. Hence, processing one task might restrict the freedom when to process another tasks. Another form of interdependency are precedence constraints between pairs of tasks. The most common precedence constraint is a minimal time lag of 0 between the completion time of the predecessor task and the start time of the successor task. In its most general from, precedence constraints impose minimal or maximal time lags on the start and/or completion times of two tasks (Neumann et al., 2003). General precedence constraints are a powerful modelling tool, which allows to depict many settings such as two tasks, which must not be processed in parallel, two tasks, which are partially processed in parallel, completion of two tasks within a given time frame, just to name a few. Closely related to precedence constraints between tasks are setup times between two tasks (Allahverdi, 2015; Hartmann and Briskorn, 2022).
Resources
There are two predominant generic concepts in the literature: machine scheduling and project scheduling. In machine scheduling resources are typically fully occupied while processing one task. Since in this case one machine does process no more than one task at a time, we often simply seek a sequence of tasks (potentially with idle time in between consecutive tasks) on each machine. Different settings with multiple machines differ in the relation of machines. In settings with parallel machines we consider alternative resources where each task requires to be processed by one machine only. These alternative resources might differ in their processing capabilities, e. g., in their speeds, or not (Pinedo, 2022). In shop settings, each task needs to be processed by each machine. The order in which machines are visited is given and identical for all tasks in a flow shop setting (Rossit et al., 2018) given and individual in a job shop and flexible job shop setting (Xiong et al., 2022; Dauzère-Pérès et al., 2024), and to be determined for each task as part of the decision process in an open shop setting (Ahmadian et al., 2021). In project scheduling resources are capable of processing multiple tasks at a time (Hartmann and Briskorn, 2010, 2022; Kolisch and Padman, 2001). Processing a task induces a certain load of capacity for each resource. As far as capacities are concerned, a set of task can be handled in parallel if the total load does not exceed the capacity of any resource. With respect to the capacity of resources there are different concepts, which can be combined. The most frequent one are renewable resources, where the capacity of a resource is available in each period anew. The capacities of non-renewable resources are available only once for all periods of the planning horizon (Weglarz et al., 2011) while the capacities of partially renewable resources are only available for subsets of the periods of the planning horizon (Alvarez-Valdes et al., 2015). In most settings capacities of resources are given and might by dynamic over time or not (Hartmann, 2015). However, there are problems where the capacity of resources has to be determined as part of the decision problem, e.g. the resource investment problem and the resource levelling problems, see Rieck et al. (2012) and Part V of Schwindt and Zimmermann (2015a).
Objectives
Beyond Deterministic Offline Scheduling
While almost all scheduling problems revolve around the components introduced above, we can distinguish different problem classes according to the nature of the information available for decision making. While in deterministic scheduling problems the parameters involved are not subject to any kind of stochasticity but specific numerical values are known, in stochastic scheduling problems one or more parameters are stochastic (Cai et al., 2014). When we encounter stochastic parameters, robustness of schedules becomes an interesting feature (Kouvelis and Yu, 1997). Robustness can be addressed with respect to feasibility of the schedule and/or the achieved objective value. In offline problems all parameters (stochastic or not) are revealed beforehand while in online problems some or maybe even all tasks or their parameters are revealed over time, (Robert and Vivien, 2009; Albers, 2009). In stochastic settings or in online settings the goal often is not to determine a full schedule immediately. Instead, a policy is determined which constructs a schedule over time when tasks or their parameters are revealed.
@@
Hierarchical planning offers a practical solution by segmenting the overall planning task into manageable subtasks (Hax and Meal, 1973), organising planning tasks into three tiers: long-term, mid-term, and short-term planning (Anthony, 1965). Long-term planning, also known as strategic planning or business planning, focuses on a multi-year horizon and shapes the company’s strategic direction, influencing decisions related to supplier selection, production capacity, and distribution structure (Stadtler, 2005). Mid-term planning comprises demand planning and master planning. It spans several months to a year and addresses aggregate-level issues such as production quantities, inventory levels, workforce requirements, and distribution channels (Pereira et al., 2020). Short-term planning, which includes scheduling and distribution planning, is conducted on a weekly or daily basis. It involves decisions regarding purchase quantities, production lot sizes, transportation modes, and vehicle routes (Bitran and Tirupati, 1993). These planning tasks are intricately linked, with a continuous flow of information across different levels, necessitating coordination to ensure alignment and integration of planning efforts across the hierarchical spectrum (Fleischmann and Meyr, 2003; Ozdamar and Yazgac, 1999). The hierarchical planning concept is illustrated in figure 2 based on Meyr et al. (2002) and Rohde and Wagner (2005).
@@
Conversely, poor PM may have severe consequences, including cost overruns, delayed timelines, and subpar quality, resulting in an organisation’s or government’s financial losses, wasted resources, and reputational damages (Bendoly and Swink, 2007; Maylor et al., 2018). High-profile project failures such as the Denver International Airport’s baggage handling system serve as reminders of the implications of poor PM. These project failures are often caused by poor planning, miscommunication or lack of collaboration between (key) stakeholders, and/or scope changes, highlighting the vital need for robust and continuous (across the project lifecycle) PM practices by all project stakeholders – individuals, groups, and/or organisations that have an interest in or are affected by the outcome of a project (Donaldson and Preston, 1995a).
The history of traditional PM tools and frameworks such as the Gantt chart and Work Breakdown Structure (Gantt, 1919; Miller, 2008) traces back to the early 20th century, with the advent of complex industrial and engineering projects necessitating more systematic approaches. However, it could be argued that early roots can be seen in the construction of the Egyptian pyramids and the Great Wall of China, where rudimentary PM principles (rather than tools) were employed (Cleland, 2004). Modern PM began to take shape during the 1950s with the development of methodologies like the Critical Path Method (CPM) and Program Evaluation and Review Technique (PERT; Malcolm et al., 1959). These methodologies formalised the way the ‘triple constraint’ of projects – time, cost, and quality – was planned and managed.

### v2

@@ Conclusion
Figure 1 gives an impression of a warehouse system consisting of a picker-to-parts and two different parts-to-picker systems: a miniload crane-based Automated Storage and Retrieval System and a system with shelf-moving robots.
Figure 1: Warehouse with a picker-to-parts and two parts-to-picker systems.
@@
Driven by technological progress, like in the fields of sensors, real-time control, autonomous driving, and robotisation, AMRs have become more and more successful in recent years (Fragapane et al., 2021). Other than their famous predecessors, Automated Guided Vehicles (AGVs), AMRs are not bound to predefined paths; computer vision allows flexible adaptation to a changing environment. In picker-to-parts systems, AMRs that also pick, such as the Toru robot of Magazino, can substitute human pickers, at least for a restricted product assortment (e.g., shoeboxes Boysen and De Koster, 2025). For non-standard products, transport AMRs accompany human pickers through the warehouse and relieve them of the burden of transporting the picked products back to the depot (Azadeh et al., 2025; Löffler et al., 2022). The most famous AMRs of parts-to-picker systems are (probably) shelf-moving (KIVA) robots. They drive under a rack (also denoted as inventory pod), lift it, and deliver the pod to a stationary picker (Weidinger et al., 2018). Other AMR representatives are storage grid-supported (Autostore) robots (Azadeh et al., 2019), rack-climbing (Exotec) robots (Chen et al., 2022), and multi-tote (Quicktron, Haipick) AMRs (Qin et al., 2024) delivering bins autonomously from storage locations. Finally, there are also AMRs for order consolidation (Zou et al., 2021). Equipped with a tiltable tray or small conveyor tablet, they can distribute goods between sorting destinations. The different decisions in picker-to-parts and parts-to-picker (including robotised) systems are summarised in figure 2.
Figure 2: Overview of decisions at different hierarchical levels in (a) picker-to-parts and (b) parts-to-picker and robotised systems.
@@
This section explores the concepts, mathematical modelling, and recent advancements in solving the BAP. For a more comprehensive description of the topic, see the latest review articles, such as Demir et al. (2014), Weiss et al. (2019), and Koyuncuoğlu (2024).
@@
BAP is an optimisation problem where the goal is to determine the optimal buffer sizes $N_{i}$ for a given set of machines, subject to constraints on system parameters. Over the years, researchers have explored a variety of objective functions and constraints.
Problem Formulation. One classical approach minimises the total buffer size while satisfying a production rate constraint:
@@
where $\bar{n}_{i}$ is the average inventory level of the $i$th buffer, and $\mathbb{E}[LT(\mathbf{N})]$ and $\mathbb{E}[WT(\mathbf{N})]$ represent the average lead time and average waiting time of customer orders. For instance, Yuzukirmizi and Smith (2008) include the lead time in their objective function.
@@
For long lines, the generalised expansion method (Kerbachea and Smith, 1987; Cruz et al., 2008) is proposed for systems with reliable machines, while decomposition and aggregation are prevailing methods to deal with unreliable machines. The decomposition method (Gershwin, 1987; Dallery et al., 1988), and its extensions (e.g., Tolio and Matta, 1998; Levantesi et al., 2003) enabled the evaluation of different models of long lines with unreliable machines. It approximates a $K$-machine $K-1$-buffer system with $K-1$ two-machine one-buffer systems. Decomposition aims to find parameters for these two-machine systems so that, for each buffer, the material flow behaviour in the two-machine line closely matches that in the original line. The goal is for an observer of the flow of parts into and out of a buffer in a real production line can be convinced that he is observing the flow of parts in a fictitious two-machine line. The aggregation method (Lim et al., 1990a; Jacobs and Meerkov, 1995; Li and Meerkov, 2007) combines all the two-machine one-buffer subsystems into a single aggregated machine. It consists of a forward aggregation which merges the first subline from the start towards the end and a backward aggregation which merges the last subline into a single machine from the end towards the beginning. This is repeated until the throughput converges. Simulation is another tool for the performance evaluation of large-scale production systems. It is flexible and can be accurate but it is time-consuming (Helber et al., 2011).
Optimisation techniques. Gradient methods (Helber, 2001; Shi and Gershwin, 2009) use decomposition to calculate gradients via forward difference, followed by a line search to locate the next optimal point. Although buffer sizes $N_{i}$s are integers, treating them as continuous variables allows gradient calculation. The optimal continuous solution is then rounded to integers (Han and Park, 2002). Meta-heuristic methods treat buffer sizes as integers. Examples include genetic algorithms (Dolgui et al., 2002), tabu search (Demir et al., 2011), simulated annealing (Spinellis et al., 2000), and degraded ceiling (Nahas et al., 2006). Hybrid algorithms that combine multiple meta-heuristics include Shi and Men (2003), Dolgui et al. (2007) and Kose and Kilincci (2015). Mathematical programming techniques include the dynamic programming which breaks the problem into smaller subproblems (Diamantidis and Papadopoulos, 2004). Colledani and Tolio (2005) linearise decomposition equations, converting the problem into a mixed-integer linear programming problem.
@@
In modern manufacturing systems, the decreasing product lifetimes necessitate frequent factory reconfigurations or replacements, limiting the time available to learn and improve these systems. As a result, intuition derived from research and from careful, systematic observation of real factories becomes increasingly vital (Gershwin, 2017). Academic research provides theoretical insights and models and offers valuable tools and methodologies that can be directly applied to improve industrial processes. Works such as Lim et al. (1990b), Burman et al. (1998), Liberopoulos and Tsarouhas (2002), Patchong et al. (2003), Colledani et al. (2010), Matta et al. (2012), and Li (2013) demonstrate how theoretical advancements in buffer allocation can be effectively applied to real-world production systems. For instance, these studies address practical challenges in industries such as electronics, food processing, automotive manufacturing, and continuous improvement in production lines, showcasing how operations research and engineering methods optimise throughput, reduce costs, and enhance system reliability.
@@
Traditionally, production control focuses on internal information flows and can generally be categorised as push (e.g., make-to-stock, driven by demand forecast) or pull (e.g., make-to-order, triggered by customer demand; Lödding, 2011; Thürer et al., 2025). With push control, the flow of materials is pushed through successive processes from suppliers to inventory and then queues, workstations, buffers, etc. until the product is complete and ready to be delivered to the customer. In pull systems, production is triggered by an order (customer demand) which pulls the materials through the processes with the aim to minimise inventory and overproduction. In modern production control systems, hybrid systems are often used to combine push and pull logic. Common hybrid approaches include drum-buffer-rope and constant work in progress (CONWIP; Spearman et al., 1990). Alternative Material Flow Control (MFC) methods have been proposed, such as order generation, order release and production authorisation (Thürer et al., 2025). The dominant trend remains with pull approaches as Lean, Kanban and Just-In-Time (JIT) systems have become established best practices for cost and waste reduction. This is further reinforced with the increased integration of external information flows with the rise of customisation, penalisation and engineer-to-order models.
@@
With Industry 4.0 developments and implementations, increased digitisation and Industrial Internet of Things (IIoT) are becoming more prevalent (Kolberg and Zühlke, 2015). However, the amount and richness of data can be overwhelming, especially for SMEs (Moeuf et al., 2018). Machine learning and predictive methods used in simulation and optimisation models are promising ways to use the data for decision making and to increase production efficiency, flexibility and resilience (Martins et al., 2020). For example, using a predictive control strategy to adapt to unpredictable changes in material inputs, such as in pharmaceutical manufacturing, can help maintain consistent product quality and yields (Eslami and Jungbauer, 2024). Digital twins for real-time monitoring and closed-loop feedback systems can address the challenges of adapting to rapidly changing manufacturing environments. For example at a process level, machine learning has been used for in-situ monitoring of metal additive manufacturing to take evasive or corrective actions when detecting avoidable defects and anomalies (Gunasegaram et al., 2024). At the manufacturing systems level, the use of real-time control methods to adapt operations to variable renewable energy supply and to support more self-sufficient and resilient manufacturing systems (Beier et al., 2017). The ISO 23247 standardisation effort on ‘Automation systems and integration – Digital twin framework for manufacturing’ are beneficial for structuring production planning and control systems (International Organization for Standardization, 2021; Shao et al., 2023), including, for example, Part 1: Overview and general principles, Part 2: Reference architecture, Part 3: Digital representation of manufacturing elements, and Part 4: Information exchange. Additional validation and more detailed specification work including validation with demonstrators are ongoing as of now. Blockchain-like technologies are also on the rise to facilitate multi-stakeholder interoperability across value chains to enable traceability and transparency in production processes and along these value chains.
@@
  • Single-model lines: A single homogeneous product is manufactured on the line (Otto et al., 2013). During each cycle time, every workstation performs the same set of tasks assigned to it. Consequently, the line-balancing problem deals with the decision how to distribute the whole set of tasks required for the product completion among the workstations.
@@
Hierarchical planning offers a practical solution by segmenting the overall planning task into manageable subtasks (Hax and Meal, 1973), organising planning tasks into three tiers: long-term, mid-term, and short-term planning (Anthony, 1965). Long-term planning, also known as strategic planning or business planning, focuses on a multi-year horizon and shapes the company’s strategic direction, influencing decisions related to supplier selection, production capacity, and distribution structure (Stadtler, 2005). Mid-term planning comprises demand planning and master planning. It spans several months to a year and addresses aggregate-level issues such as production quantities, inventory levels, workforce requirements, and distribution channels (Pereira et al., 2020). Short-term planning, which includes scheduling and distribution planning, is conducted on a weekly or daily basis. It involves decisions regarding purchase quantities, production lot sizes, transportation modes, and vehicle routes (Bitran and Tirupati, 1993, see also §2.15). These planning tasks are intricately linked, with a continuous flow of information across different levels, necessitating coordination to ensure alignment and integration of planning efforts across the hierarchical spectrum (Fleischmann and Meyr, 2003; Ozdamar and Yazgac, 1999). The hierarchical planning concept is illustrated in figure 4 based on Meyr et al. (2002) and Rohde and Wagner (2005).
@@
Planning problems are often complex due to the interplay of multiple objectives, constraints, time horizons, and decision layers. To address these challenges, numerous mathematical modelling approaches have been developed, ranging from mixed-integer (linear) programming (e.g., Bilgen and Ozkarahan, 2007; Dogan and Goetschalckx, 1999) to fuzzy programming (e.g., Peidro et al., 2010), or stochastic programming (e.g., Sabri and Beamon, 2000), aiming to obtain optimal or near-optimal solutions. The richness of this modelling landscape reflects the diversity of planning environments, each requiring tailored formulations to capture domain-specific characteristics. Reviews by Fleckenstein et al. (2023), Gelders and Van Wassenhove (1981), and Mula et al. (2010) provide valuable overviews of optimisation-based models that support planning.
Scheduling is the allocation of resources to tasks over time. The result is a schedule, which gives information on the start time and the completion time of each task as well as the times when resources are processing (which) tasks throughout the planning horizon. Synonyms for tasks in the literature are jobs or activities. Scheduling is one of the major and classical planning problems in operations management, see, e.g., §2.4, §2.11, §2.16, §2.18. Recently, scheduling is also addressed in supply chain management, see Amaro and Barbosa-Póvoa (2008), Chen and Hall (2022), Kreipl and Pinedo (2004), and Sawik (2011). In most cases, scheduling is a short- to mid-term planning problem, where the planning horizon stretches from a few hours to several months. Scheduling is done in manufacturing and services, see, e.g., Pinedo (2009). In manufacturing, tasks relate to physical products, which have to be processed in production facilities. In services, the main fields for scheduling problems are in transportation (e.g., trucks, which have to be (un)loaded at terminal gates, see Boysen and Fliedner, 2010, ships which have to be (un)loaded at berths, see Bierwirth and Meisel, 2010, and Bierwirth and Meisel, 2015, containers, which have to be moved by cranes, see Boysen et al., 2017, aircrafts, which have to use runways for starting and landing, see Ikli et al., 2021), healthcare (e.g, patients, which have to be treated by doctors, see Hall, 2012, and Youn et al., 2022), sports (see Kendall et al., 2010; Ribeiro, 2013)), and workforce scheduling (see Van den Bergh et al., 2013). See also Pinedo (2009) and Pinedo et al. (2015) for a survey of scheduling in services.
Tasks
While being processed a task occupies resources. In the most simple case the resource is occupied completely and cannot handle any other tasks in parallel. However, tasks also might seize a fraction of a resource’s capacity and, if it is sufficiently small, the resource might handle other tasks in parallel. A task’s processing time specifies how long it occupies the resource. In most cases, the processing time is a given parameter. In more involved settings the decision maker might have the opportunity to decide on the processing time (Weglarz et al., 2011; Shabtay and Steiner, 2007; Gawiejnowicz, 2020). Once the processing of a task has started it may or may not be preempted. However, processing a task is only allowed within a time interval between its release date and its deadline. In (quite common) special cases, release dates and deadlines may not impose any relevant restrictions and we can think of them as non-existent.
Precedence constraints
The times when tasks are processed are interdependent through the joint need for scarce resources. Hence, processing one task might restrict the freedom when to process another tasks. Another form of interdependency is precedence constraint between task pairs. The most common precedence constraint is a minimal time lag of 0 between the completion time of the predecessor task and the start time of the successor task. In its most general from, precedence constraints impose minimal or maximal time lags on the start and/or completion times of two tasks (Neumann et al., 2003). General precedence constraints are powerful modelling tools, allowing to depict many settings such as two tasks, which must not be processed in parallel or are partially processed in parallel, or completion of two tasks within a given time frame, just to name a couple (Hartmann and Briskorn, 2022).
Resources
There are two predominant generic concepts in the literature: machine scheduling and project scheduling. In machine scheduling resources are typically fully occupied while processing one task. Since in this case one machine does process no more than one task at a time, we often simply seek a sequence of tasks (potentially with idle time in between consecutive tasks) on each machine. Different settings with multiple machines differ in the relation of machines. In settings with parallel machines we consider alternative resources where each task requires to be processed by one machine only. These alternative resources might differ in their processing capabilities, e. g., in their speeds, or not (Pinedo, 2022). In shop settings, each task needs to be processed by each machine. The order in which machines are visited is given and identical for all tasks in a flow shop setting (Rossit et al., 2018) given and individual in a job shop and flexible job shop setting (Xiong et al., 2022; Dauzère-Pérès et al., 2024), and to be determined for each task as part of the decision process in an open shop setting (Ahmadian et al., 2021). In project scheduling resources are capable of processing multiple tasks at a time (Hartmann and Briskorn, 2010, 2022; Kolisch and Padman, 2001). Processing a task induces a certain load of capacity for each resource. As far as capacities are concerned, a set of tasks can be handled in parallel if the total load does not exceed the capacity of any resource. With respect to the capacity of resources there are different concepts, which can be combined. The most frequent one are renewable resources, where the capacity of a resource is available in each period anew. The capacities of non-renewable resources are available only once for all periods of the planning horizon (Weglarz et al., 2011) while the capacities of partially renewable resources are only available for subsets of the periods of the planning horizon (Alvarez-Valdes et al., 2015). In most settings capacities of resources are given and might by dynamic over time or not (Hartmann, 2015). However, there are problems where the capacity of resources has to be determined as part of the decision problem, e.g. the resource investment problem and the resource levelling problems, see Rieck et al. (2012) and Part V of Schwindt and Zimmermann (2015a).
Objectives
Solution methods
Scheduling problems are combinatorial in nature and the vast majority is ${\cal NP}$-hard. Hence, classical methods for solving scheduling problems exactly are mixed-integer linear programs (MIPs), branch-and-bound, and dynamic programming (Agnetis et al., 2025). Constraint programming (CP; Baptiste et al., 2012) is a solution approach, which is particularly suited to scheduling problems. Recent results show that for certain scheduling problems, CP is superior to MIPs (Naderi et al., 2023). Additionally, combining MIP and CP is a promising solution strategy (Pohl et al., 2022). Heuristics such as priority rules and metaheuristics are used to solve large problem instances (Agnetis et al., 2025; Ruiz, 2018). Furthermore, matheuristics and CP heuristics successfully combine heuristic techniques with MIP- and CP-approaches (Agnetis et al., 2025). Recent research investigates machine learning methods, most prominently reinforcement learning and deep reinforcement learning (Kayhan and Yildiz, 2023; Khadivi et al., 2025). For the latter, graph neural networks are of particular interest, since the precedence constraints between tasks are naturally represented by graphs (Smit et al., 2025).
Beyond deterministic offline scheduling
While almost all scheduling problems revolve around the components introduced above, we can distinguish different problem classes according to the nature of the information available for decision making. While in deterministic scheduling problems the parameters involved are not subject to any kind of stochasticity but specific numerical values are known, in stochastic scheduling problems one or more parameters are stochastic (Cai et al., 2014). When we encounter stochastic parameters, robustness of schedules becomes an interesting feature (Kouvelis and Yu, 1997). Robustness can be addressed with respect to feasibility of the schedule and/or the achieved objective value. In offline problems all parameters (stochastic or not) are revealed beforehand while in online problems some or maybe even all tasks or their parameters are revealed over time (Robert and Vivien, 2009; Albers, 2009). In stochastic settings or in online settings the goal often is not to determine a full schedule immediately. Instead, a policy is determined which constructs a schedule over time when tasks or their parameters are revealed.
@@
Conversely, poor PM may have severe consequences, including cost overruns, delayed timelines, and subpar quality, resulting in an organisation’s or government’s financial losses, wasted resources, and reputational damages (Bendoly and Swink, 2007; Maylor et al., 2018). High-profile project failures such as the Denver International Airport’s baggage handling system serve as reminders of the implications of poor PM. These project failures are often caused by poor planning, miscommunication or lack of collaboration between (key) stakeholders, and/or scope changes, highlighting the vital need for robust and continuous (across the project lifecycle) PM practices by all project stakeholders – individuals, groups, and/or organisations that have an interest in or are affected by the outcome of a project (Donaldson and Preston, 1995).
The history of traditional PM tools and frameworks such as the Gantt chart and Work Breakdown Structure (Gantt, 1919; Miller, 2008) traces back to the early 20th century, with the advent of complex industrial and engineering projects necessitating more systematic approaches. However, it could be argued that early roots can be seen in the construction of the Egyptian pyramids and the Great Wall of China, where rudimentary PM principles (rather than tools) were employed (Cleland, 2004). Modern PM began to take shape during the 1950s with the development of methodologies like the Critical Path Method (CPM; Meredith et al., 2021) and Program Evaluation and Review Technique (PERT; Malcolm et al., 1959). These methodologies formalised the way the ‘triple constraint’ of projects – time, cost, and quality – was planned and managed.